OpenStudy (anonymous):

Choose the missing number to make equivalent fractions. 1/4 = /16 A. 2 B. 4 C. 6 D. 8

OpenStudy (anonymous):

Are you adding or multiplying?

OpenStudy (anonymous):

you have to find the missing number in: /16 to make a equvilent fraction

OpenStudy (anonymous):

Right. I know that. But it would help to find the number if you you knew to add or multiply. Did your textbook not tell you, or can you do either one?

OpenStudy (anonymous):

It didn't tell me

OpenStudy (anonymous):

\[\frac{1}{4}=\frac{x}{16}\] Solve for x (multiply both sides by 16).

OpenStudy (anonymous):

i think its B or C

OpenStudy (anonymous):

You can multiply both sides by 16 to isolate x. What's \[\frac{1}{4}*16\]?

OpenStudy (anonymous):

1/4 x 16/1 = 16/4- Simplify-1 whole?

OpenStudy (anonymous):

k i got the answer

OpenStudy (anonymous):

Okay! :)